OpenStudy (anonymous):

what is the remainder of 3x^4+2x^3-x^2+2x-19 divided by x+2

OpenStudy (amistre64):

what is the value of the long part when x+2 = 0?

OpenStudy (amistre64):

\[P(x)=(x-a)Q(x)+R(x)\] \[P(a)=(a-a)Q(x)+R(a)\] \[P(a)=0+R(a)\]

OpenStudy (anonymous):

m just doing long division but I got a brain fart

OpenStudy (amistre64):

hmm, show mw what youve worked out so far, i can prolly see where you might be having an issue at

OpenStudy (amistre64):

\[\frac{P(x)}{(x-a)}=Q(x)+\frac{R(x)}{x-a}\] if you just need the R(x), then working out the Q(x) is just alot of extra work